# IKUN for WMT24 General MT Task: LLMs Are here for Multilingual Machine Translation

### v1

@@ 3 Tokenizer Efficiency
For our open system, IKUN, we didn’t modify its tokenizer, since Llama-3 tokenizer is already efficient enough, only except for Hindi. For our constrained system, IKUN-C, we expanded its vocabulary for Chinese, Japanese, Hindi, and Icelandic through the following steps: (1) Generate a new vocabulary of 120K sub-words using monolingual data from the past two years for these four languages from News Crawl [^3] https://data.statmt.org/news-crawl/; and (2) Merge the new vocabulary with the original. The efficiency of the resulted IKUN-C tokenizer is shown in Figure 1, demonstrating more efficiency for these four languages, especially for Hindi.

### v2

@@ 3 Tokenizer Efficiency
For our open system, IKUN, we didn’t modify its tokenizer, since Llama-3 tokenizer is already efficient enough, only except for Hindi. For our constrained system, IKUN-C, we expanded its vocabulary for Chinese, Japanese, Hindi, and Icelandic through the following steps: (1) Generate a new vocabulary of 12K sub-words using monolingual data from the past two years for these four languages from News Crawl [^3] https://data.statmt.org/news-crawl/; and (2) Merge the new vocabulary with the original. The efficiency of the resulted IKUN-C tokenizer is shown in Figure 1, demonstrating more efficiency for these four languages, especially for Hindi.